# Generative Gaussian Splatting for Unbounded 3D City Generation

### v2

@@ 2 Related Work
Scene Generation. Unlike impressive models for generating objects and avatars, generating scenes presents greater challenges due to their extremely high diversity. Earlier approaches [27, 33] generate scenes by synthesizing videos, which lack 3D awareness and cannot ensure 3D consistency. Semantic image synthesis methods [17, 40, 49] exhibit promising results in generating scene-level content by conditioning on pixel-wise dense correspondence, such as semantic segmentation maps or depth maps. Several methods [6, 11, 59] generate natural 3D scenes by performing inpainting and outpainting on RGB images or feature maps. However, most methods can only interpolate or extrapolate a limited distance from the input views and lack generative capabilities. Recent approaches [10, 31, 55] have achieved 3D consistent scenes at an infinite scale through unbounded layout extrapolation. Another set of works [3, 13, 41] focuses on indoor scene synthesis using costly 3D datasets [12] or CAD object retrieval [14, 15].
@@ 3 Our Approach
Figure 2: The framework of GaussianCity. To create an unbounded 3D city, the BEV points are firstly generated from a local patch of the BEV maps, which includes the height field $\mathbf{H}$, semantic map $\mathbf{S}$, and binary density map $\mathbf{D}$. Then, the BEV-Point attributes $\left\{\mathbf{I},\mathbf{C}_{\it A},\mathbf{C}_{\it R},\mathbf{F}_{\it S}\right\}$ are generated for each point and the Style Lookup Table: $\mathcal{T}(\mathbf{L})\to\mathbf{Z}_{T}$ is generated for each instance. Next, the BEV-Point decoder generates the Gaussian attributes $\mathbf{A}$ from BEV-Point attributes. Finally, the Gaussian Rasterizer $\mathcal{R}$ produces the rendered image $\mathbf{R}$.
@@ 5 Conclusion
Acknowledgements. This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).

### v3

@@ 2 Related Work
Scene Generation. Unlike impressive models for generating objects and avatars, generating scenes presents greater challenges due to their extremely high diversity. Earlier approaches [26, 32] generate scenes by synthesizing videos, which lack 3D awareness and cannot ensure 3D consistency. Semantic image synthesis methods [16, 40, 48] exhibit promising results in generating scene-level content by conditioning on pixel-wise dense correspondence, such as semantic segmentation maps or depth maps. Several methods [7, 13, 59] generate natural 3D scenes by performing inpainting and outpainting on RGB images or feature maps. However, most methods can only interpolate or extrapolate a limited distance from the input views and lack generative capabilities. Recent approaches [11, 30, 54, 4] have achieved 3D consistent scenes at an infinite scale through unbounded layout extrapolation. Another set of works focuses on indoor scene synthesis using costly 3D datasets [3, 14] or CAD object retrieval [41].
@@ 3 Our Approach
Figure 2: The framework of GaussianCity. To create an unbounded 3D city, the BEV points are firstly generated from a local patch of the BEV maps, which includes the height field $\mathbf{H}$, semantic map $\mathbf{S}$, and binary density map $\mathbf{D}$. Then, the BEV-Point attributes $\left\{\mathbf{I},\mathbf{C}_{\it A},\mathbf{C}_{\it R},\mathbf{F}_{\it S}\right\}$ are generated for each point and the Style Lookup Table: $\mathcal{T}(\mathbf{L})\to\mathbf{Z}_{T}$ is generated for each instance. Next, BEV-Point Decoder generates the Gaussian attributes $\mathbf{A}$ from BEV-Point attributes. Finally, Gaussian Rasterizer $\mathcal{R}$ produces the rendered image $\mathbf{R}$.
@@ 5 Conclusion
Acknowledgements. This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221- 0012, MOE-T2EP20223-0002), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).